OpenStudy (eric_d):

Domain of function h , where h(x)2^x, is real number a)find images of 0,2,-1 and 6 @ganeshie8

OpenStudy (eric_d):

Do I need to substitute the given values to the function

ganeshie8 (ganeshie8):

Yes !

OpenStudy (eric_d):

Okay

ganeshie8 (ganeshie8):

h(x) = 2^x you just need to find h(0), h(2), h(-1), etc..

ganeshie8 (ganeshie8):

http://en.wikipedia.org/wiki/Image_(mathematics)

OpenStudy (eric_d):

Understood Another question IF p(X)=2X^3+4X^2-X+3 AND q(X)=3X^2+X-1 FIND A) P(x) times Q(x)

ganeshie8 (ganeshie8):

multiply

ganeshie8 (ganeshie8):

\[P(x) \times Q(x) = (2x^3+4x^2-x+3 )\times \color{red}{(3x^2+x-1 )}\]

OpenStudy (eric_d):

I'm nt sure bout the factorisation part Can u show me the first step.. I can do the rest

ganeshie8 (ganeshie8):

we need to distribute each term of one factor with the other factor

ganeshie8 (ganeshie8):

\[\large P(x) \times Q(x) = (2x^3+4x^2-x+3 )\times \color{red}{(3x^2+x-1 )}\] \[\large = (2x^3+4x^2-x+3 )\times \color{red}{(3x^2)} \\ \large + (2x^3+4x^2-x+3 )\times \color{red}{(x)} \\ \large + (2x^3+4x^2-x+3 )\times \color{red}{(-1)} \]

OpenStudy (eric_d):

Now, I undersatand the concept Thanks a lot

OpenStudy (eric_d):

understand*

ganeshie8 (ganeshie8):

good :) its going to be a bit lengthy - after multiplying, you need to group the like terms and simplify

OpenStudy (eric_d):

Okay Another Q: Find quotient if x^3-4x^2+5x-2 can be divided by (x-2) exactly

OpenStudy (eric_d):

this is a sample Q Hang on

ganeshie8 (ganeshie8):

okie

OpenStudy (eric_d):

the solution is x^3-4x^2+5x-2=q(x) . (x-2) They mention that q(x) is a polynomial of degree 2.. why?

ganeshie8 (ganeshie8):

whats the degree of polynomial on LEFT HAND side ?

OpenStudy (eric_d):

3

ganeshie8 (ganeshie8):

what happens to its degree if you multiply the entire polynomial by \(x\) ?

OpenStudy (eric_d):

It'll be 4.. I'm nt sure

ganeshie8 (ganeshie8):

Yes ! multiplying x^3 by x gives u x^4 so clearly the degree is 4

OpenStudy (eric_d):

okay

ganeshie8 (ganeshie8):

Now look at the question you have asked earlier

ganeshie8 (ganeshie8):

x^3-4x^2+5x-2=q(x) . (x-2) whats the degree of q(x) ?

ganeshie8 (ganeshie8):

Since the degree of LEFT SIDE polynomial is `3`, the degree ofRIGHT SIDE polynomial must also equal `3`

ganeshie8 (ganeshie8):

Since `(x-2)` has degree `1` , the degree of q(x) has to be `2` so that when you multiply both you get a polynomial of degree 2+1 = 3

OpenStudy (eric_d):

Understood

OpenStudy (eric_d):

One more question Using remainder theorem, find the remainder when polynomial P(x)=2x^3+7x^2-5x-4 is divided by (x+3)

OpenStudy (eric_d):

do i jst need to find the x-intercept

ganeshie8 (ganeshie8):

As the question says, you need to divide and find the remainder

OpenStudy (eric_d):

Nvm, I gt it x+3=0 x=-3

OpenStudy (eric_d):

Thank you Ganesh

ganeshie8 (ganeshie8):

Using remainder theorem, find the remainder when polynomial P(x)=2x^3+7x^2-5x-4 is divided by (x+3)

ganeshie8 (ganeshie8):

x=-3 evaluate the polynomial at -3 P(-3) = ?

OpenStudy (eric_d):

=2(-3)^3+7(-3)^2-5(-3)-4 =

OpenStudy (eric_d):

20

OpenStudy (eric_d):

Remainder is 20

ganeshie8 (ganeshie8):

Excellent !!

OpenStudy (eric_d):

Alrite, thanks again I'm done 4 nw... :)